OpenStudy (anonymous):

I will give medals!! Please help!!!! Let f(x)=-2x+4 and g(x)=-6x-7. Find f(x)-g(x).

OpenStudy (anonymous):

here is how you set up the problem.... f(x)-g(x) -2x+4-(-6x-7)... Make sure to distribute the negative... -2x+4+6x+7 then combine like terms

OpenStudy (anonymous):

f(x) =-2x +4 g(x) = -6x -7 subtract them, term by term, -2x-(-6x) = -2x+6x = 4x etc.

OpenStudy (anonymous):

@Gerardo_cast23 So you pretty much just plug in the f(x) and the g(x), then simplify???